OpenStudy (anonymous):

The length of a rectangle is 1ft less than twice its width, and the area of the rectangle is 78ft^2. Find the dimensions of the rectangle.

Directrix (directrix):

Let w = width Length is 2w - 1 78 = w (2w - 1) 2w^2 - w - 78 = 0 (w+6)(2w-13) = 0 w = -6 or 2w = 13 w = -6 or w = 13/2 -6 does not apply so w = 6.5 Dimensions of the rectangle are 6.5 feet by 12 feet ---> Ans

OpenStudy (anonymous):

Thank you...

Directrix (directrix):

Glad to help.